ISOLDE Technical Coordination Meeting (LS3)

Europe/Zurich
508/1-001 (CERN)

508/1-001

CERN

20
Show room on map
Zoom Meeting ID
62093336977
Host
Joachim Vollaire
Useful links
Join via phone
Zoom URL
    • 14:00 14:05
      General news - Review of minutes 5m
      Speakers: Erwin Siesling (CERN), Joachim Vollaire (CERN)
    • 14:05 14:15
      Safety matters 10m
      • See attachment for directional traffic outside b.179 / ISOLDE parking lot
    • 14:15 14:20
      IBDRS project - activities update 5m
      Speakers: Ana-Paula Bernardes (CERN), Joachim Vollaire (CERN)
      • Excavation ongoing. Up to 25 trucks per day
      • Core sampling ongoing.
      • New plan for the vehicle traffic (see safety point)
    • 14:20 14:25
      Ventilation primary areas upgrade 5m

      Activities to be coordinated with platform installation:

      1. EN-CV RW disposal - Planned the week of the 2 Feb (exact day tbc)
      2. Equipment transfer from 170/3-402 to 197/2-401- 12Feb
      3. Fire detection control rack installation on the platform - 18 Feb
      Speakers: Joachim Vollaire (CERN), Melania Averna (CERN)
    • 14:25 14:30
      RILIS Laboratory upgrade project - activities update 5m
      Speakers: Carlotta Porzio (CERN), Katerina Chrysalidis (CERN), Laura Lambert (CERN), Sebastian Rothe (CERN)
      • Cables
        • For suspending the cables SCE has said they can easily attach them onto the scaffolding - however, this would mean they would still require suspension for 1 week as the time it takes to detach the cables, then build scaffolding, then attach cables to scaffolding this would be the estimated lead time.
        • A new on-site visit will be requested in order to see if a temporary frame can be built over the beamline in order to rest the cables on that, but there is limited space
          • If possible / space allows, mechanical protection over the beamline is preferabe as this would then protect even from the construction of the scaffolding itself
      • Protection of equipment
        • The idea is this will now be done by RBS. Timeline still TBD.
      • Working footprint attached
        • This shows the area that will be worked on taking into account both the dismantling as well as the construction phases
      • GPS exhaust vac pipes
        • Re-routing not possible
        • plan b - will design a permanent shielded box
      • Dismantling
        • CV
          • require 2 weeks
          • space for 2 pallets for dismantling of AC units and ducts
          • Route for removal of pallets?
        • ACE
          • 3 weeks end of may/ early june
          • green cage and light structure
          • Route to be defined and location of pallets to be defined
      • Integration
        • Meeting tomorrow afternoon for revision of current integration
      • Planning
        • Carlotta to come up with a master plan which integrates all groups on a weekly block schedule
        • This can then be circulated among relevant groups by management
        • ECR for EP cage change?
    • 14:45 14:50
      Activities in the primary areas - Coactivity management 5m
      Speaker: Joachim Vollaire (CERN)
    • 14:50 14:55
      Activities in the expeirmental hall - Coactivity management 5m
      Speakers: Erwin Siesling (CERN), Hanne Heylen (CERN), Oscar Fjeld (CERN)
      • Refurbishment of the power outlets ongoing (EN-EL / James Devine). This will last the whole week. Precautions will be taken no to cut the power for the WISARD setup (ongoing tests this and next week).
      • The overhead crane remote is back in the cuboard. PR214 overhead crane can be used by those who have the license again.
      • Re-zoning of the overhead crane PR214 in the WISARD area by the transport team (Yann Seraphin) and dewar handling to WISARD on Tuesday 27 Jan.
         
      • EPC platform:
        - Construction work of the EPC platform (508 side) ongoing. Crane blockage foreseen next week 4-5(-6) Feb for the dismounting of the scaffolding.
        - Helmets have been moved to the left after entrance to the 1st floor of 197. Final position will be on the new cage of the EPC platform once the installation has finished.
        - Restricted access from the entrance to the hall from the 508 side during the EPC platform construction work. Visits to the hall from the SAS side (helmets available). Estimate until 6 Feb.
      • Request Rustan Calming to store 16 racks for the EPC platform already to the hall: Asked to wait until installation starts (no space in the hall). (Erwin)
      • Dust from the construction work is provoking false fire alarms: the fire brigade asked to keep the doors to 197 and the old Control Room closed. Signs have been put in place. Reminder: The key for the door towards 197-1 power room is SC40. Nr 33 in the ISOLDE TRAKA cupboard.
         
      • Cable activities in the hall: 
        - New power cables distribution for the CHROME RP detectors. New distribution box on the 170 hall wall at the Salve side next to the RP buffer zone. Cables to the HEBT lines and tunnel side (routing through the HIE mezzanine). Partial cabling to the RILIS zone (final installation after RILIS work finishes).
        - Power cable for warming gas bottles (new flamable gas rack outside 170, Rte Einstein side) foreseen in parallel.
        Contact: Jerome Cantet EN-EL-SW

      • Recuperation bins for metal, etc. presently in front of B198 (Cryo) to be moved in front of the (to be) diesel position at B199. With the restricted and difficult access to the IBDRS side this is a more logic and useful position. Thierry Dupont (Cryo TSO 198/199) has been informed and is OK (Hanne).

       

      • 508 access road:
        • Installation of concrete slab for MULTIPAC - need to close access road for 4 days - VIC tomorrow 28/1, asked to do this Thursday 29/1 and 30/1 -> I will give feedback based on this meeting
        • remark Erwin: Would this not clash with the recuperation of the old ventilation unit material bins next to 197 as well as with the construction work on the EPC platform (contractors van / material)?

    • 14:55 15:00
      Planned visits 5m
      Speaker: Hanne Heylen (CERN)
      • Request to move the Technical Meeting to B37 due to a protocol office visit by the Maastricht University foreseen Tuesday 3 Feb at 14h.
      • List until Tues 10/2